OpenStudy (anonymous):

Find the force of gravity between a newborn baby (mass = 2.9 kg) and the planet Mars (mass = 6.4 x 10^23kg), when Mars is at its closest to Earth (distance = 5.6 x 10^10m).

OpenStudy (anonymous):

ok this is simple

OpenStudy (anonymous):

\[F = u \frac{mM}{r^2}\]

OpenStudy (anonymous):

small m represent the mass of the baby, M is the mass of Mars

OpenStudy (anonymous):

\[F = u \frac{2.9kg \times 6.4 \times 10^{23} kg}{(5.6 \times 10^{10}m^2)}\]

OpenStudy (anonymous):

\[F=6.67300 × 10^{-11} m^3 kg^{-1} s^{-2} \frac{2.9kg \times 6.4 \times 10^{23} kg}{(5.6 \times 10^{10}m^2)}\]

OpenStudy (anonymous):

m/s^2 is = N

OpenStudy (anonymous):

SO how would I write it as an answer.

OpenStudy (anonymous):

all u have to do is use the calculator

OpenStudy (anonymous):

\[F=6.67300 × 10^{-11} \frac{2.9 \times 6.4 \times 10^{23} }{(5.6 \times 10^{10})}\]

OpenStudy (anonymous):

is that clearer for u now?

OpenStudy (anonymous):

do u know how to use a calculator?

OpenStudy (anonymous):

To be honest with you, I don't know how to use the Texas Instrument calculator. I'm 41 and we didn't have these in my time and if we did we couldn't afford one. My kids are the lucky ones who got to use it. But they are no help.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

I answered it. 3.9 x 10^-8 N